An electrically linear polar molecule has positive charge equivalent to the charge on one proton centre's 0.30 nm from its center of mass and an equal quantity of negative charge centered 0.10nm from its center of mass{opposite the positive charge}. The molecule is within an electric field of 120 N/C.

What it's the maximum torque experienced by the molecule?

What is the minimum torque experienced by the molecule
1
Expert's answer
2021-09-02T10:41:34-0400

Given:

"|q_1|=|q_2|=e=1.6*10^{-19}\:\rm C"

"d=d_1+d_2=0.30\:\rm nm+0.10\:\rm nm=0.40\:\rm nm"

"E=120\:\rm N/C"


The electric dipole moment is given by

"p=qd=1.6*10^{-19}*0.40*10^{-9}=6.4^{-29}\:\rm C\cdot m"

The torque experienced by the molecule

"\bf \tau=p\times E"

The maximum torque

"\tau=pE=6.4^{-29}*120=7.7*10^{-27}\:\rm N\cdot m"

The minimum torque

"\tau=0"